# Alternating Minimization Schemes for Computing Rate-Distortion-Perception Functions with $f$-Divergence Perception Constraints

### v1

@@ IV-A Asymptotic Convergence Rate Analysis
We remark that the iteration structure in Theorem 2, i.e., $h^{(n+1)}$ = $q_{Y}[h^{(n)}]$ given the current iteration $n$, can be described as an implicit vector function $S:\mathbb{R}^{|\mathcal{X}|}\to\mathbb{R}^{|\mathcal{X}|}$ with $S[h](i)=h(i)\cdot c[h,S[h]](i)$, such that $h^{(n+1)}=S[h^{(n)}]$. The results of Lemma 2 characterize a distribution $h^{*}$ achieving the RDPF as a fixed point of $S(h)$, i.e. $h^{*}=S[h^{*}]$, since $c[h^{*},S[h^{*}]](i)=1,~{}i=1\ldots,|\mathcal{X}|$. Under these observations, we can analyze the convergence rate of the OAM scheme following similar steps as in [42].
The first order Taylor expansion of $S[h]$ around a fixed point $h^{*}$ is defined as
$S[h]=S[h^{*}]+J[h^{*}]\cdot(h-h^{*})+o(||h-h^{*}||)$
@@
(Jacobian form) The Jacobian $J(h)$ computed at the fixed point $h^{*}$ is given as
@@
where $M^{*}=M[h^{*},h^{*}]$ and $\Gamma^{*}=\Gamma[h^{*},h^{*}]$ as defined in (LABEL:matrix:_Newton_M) and (14), respectively.
@@
Let $\{\lambda_{i}\}_{i\in[1:|\mathcal{X}|]}$ be the set of eigenvalues of $M^{*}=M[h^{*},h^{*}]$. Given a distortion function $d:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{0}^{+}$ that induces a full-rank matrix $D=[e^{-s_{D}d(i,j)}]_{(i,j)\in\mathcal{X}^{2}}$, then $\lambda_{i}>0$,$\forall i\in[1:|\mathcal{X}|]$, i.e., $M^{*}$ has only positive eigenvalues.
@@
$1\displaystyle 0\leq\{\theta_{i}\}_{i\in[1:|\mathcal{X}|]}<$ 1 .
Moreover, let $\gamma\in[\theta_{\max},1)$. Then, there exists $\delta>0$ and $K>0$ such that if $h^{(0)}\in\{h:||h-h^{*}||\leq\delta\}$, we obtain
$||h^{(n)}-h^{*}||<K\cdot||h^{(0)}-h^{*}||\cdot\gamma^{n}$ (21)
@@
The convergence rate of the NAM scheme is immediate from the OAM scheme analysis, given the close relation between the two schemes. Since the only difference lies in the introduction of Newton’s root-finding method for the estimation of the optimal iteration step, the NAM scheme enjoys the same convergence rate in terms of the number of iterations as the OAM scheme, i.e., an exponential convergence $\mathcal{O}(e^{-n})$ under the assumptions of Lemma 5. However, the added complexity from the application of Newton’s method at each iteration increases the overall iteration complexity, due to the at least linear convergence rate $\mathcal{O}(\frac{1}{m})$ of the root approximant. Therefore, the total complexity is approximately $\mathcal{O}(\frac{e^{-n}}{m})$, where $n$ and $m$ depend on the error tolerances $\epsilon$ and $\delta$ given as input in Algo. 1.
@@
where $M$ and $\Gamma$ are given by (LABEL:matrix:_Newton_M) and (14), respectively. Unlike Theorem 2, where the structure of (20) bounds its own eigenvalues, in this case, we need to bound the Lagrangian multiplier $s_{P}$, hence matrix $\Gamma$, to guarantee exponential convergence of the algorithm. This is proved in the following theorem.
@@
For a given $s_{D}\geq{0}$, let $I_{s_{P}}=[0,s_{P,\max}]$ be the domain of $s_{P}$, $\{\theta_{a,i}\}_{i\in\mathcal{X}}$ the set of eigenvalues of $\hat{J}(h^{*})$ and $\theta_{\max}$ the maximum eigenvalue of $J(h^{*})$ in (20). Define the set $I_{s_{P}}^{\epsilon}=[0,s_{P,\max}-\epsilon]$ for $0<\epsilon<s_{P,\max}$. Then, there exists an $\epsilon^{\prime}$ such that if $s_{P}\in I_{s_{P}}^{\epsilon^{\prime}}$ then $10\leq\{\theta_{a,i}\}_{i\in\mathcal{X}}<$ 1.
@@ V Numerical Results
In this section, we validate our theoretical findings with simulations. In particular, we consider the computation of the RDPF under the NAM and RAM schemes, respectively using Algorithms 1 and 2.
@@ V-A RDPF Computation - NAM scheme
Suppose that ${\cal X}=\{0,1\}$ with $p_{X}\sim Ber(0.15)$, and let $d(\cdot,\cdot)=d_{H}(\cdot,\cdot)$ with perception constraint chosen to be either (a) $D_{f}(\cdot||\cdot)=D_{JS}(\cdot||\cdot)$, (b) $D_{f}(\cdot||\cdot)=D_{KL}(\cdot||\cdot)$, (c) $D_{f}(\cdot||\cdot)=D_{\chi^{2}}(\cdot||\cdot)$, (d) $D_{f}(\cdot||\cdot)=D_{\alpha=-1}(\cdot||\cdot)$, (e) $D_{f}(\cdot||\cdot)=D_{\alpha=\frac{1}{4}}(\cdot||\cdot)$, (f) $D_{f}(\cdot||\cdot)=D_{\alpha=\frac{1}{2}}(\cdot||\cdot)$. In Fig. 1, we display the $R(D,P)$ estimates resulting from Algorithm 1 for each divergence metric.
@@
(e) $D_{\alpha=\frac{1}{4}}$
(f) $D_{\alpha=\frac{1}{2}}$
Figure 1: $R(D,P)$ for a Bernoulli source under a Hamming distortion and various perception constraints.
We observe that all the computed RDPFs share similarities in the structure of the operating regions on the $(D,P)$ plane. Taking as a reference the plot of Fig. 1(a), we distinguish three cases:
  • Case I, where the perception level $P$ is not met with equality. In this region, the RDPF is equal to the RDF.
@@
The results show that Algorithm 1 can completely cover the Case III region. The boundary between Cases I-III (black line) representing the RDF is obtained by setting the Lagrangian multiplies $s_{P}=0$. Furthermore, the Case I region can be obtained by extension of the relative boundaries, since all the $(D,P)$ points share the solution given by the RD problem.
@@ V-B RDPF Computation - RAM scheme
Suppose that ${\cal X}=\hat{\cal X}=\{0,1\}$ with $p_{X}\sim Ber(0.15)$, and let $d(\cdot,\cdot)=d_{H}(\cdot,\cdot)$ with perception constraint chosen to be either (a) $D_{f}(p_{X}||h)=D_{JS}(\cdot||\cdot)$, (b) $D_{f}(\cdot||\cdot)=D_{KL}(\cdot||\cdot)$, (c) $D_{f}(\cdot||\cdot)=D_{\chi^{2}}(\cdot||\cdot)$, (d) $D_{f}(\cdot||\cdot)=D_{\alpha=-1}(\cdot||\cdot)$, (e) $D_{f}(\cdot||\cdot)=D_{\alpha=\frac{1}{2}}(\cdot||\cdot)$, (f) $D_{f}(\cdot||\cdot)=TV(\cdot||\cdot)$. In Fig. 2, we display the $R(D,P)$ estimates resulting from Algorithm 2 for each divergence metric.
@@
(e) $D_{\alpha=\frac{1}{2}}(\cdot||\cdot)$
(f) $TV(\cdot||\cdot)$
Figure 2: $R(D,P)$ for a Bernoulli source under a Hamming distortion and various perception constraints.
Remarks similar to those in Example 1 can be made regarding the operating regions of the computed RDPFs. In addition, in Fig. 2(f) we compare the theoretical results of [3, Equation 6] with the numerical results obtained using Algorithm 2. This case is of particular interest due to the non-differentiability of the $TV$ distance. We observe that Algorithm 2 achieves exactly the theoretical solution of [3, Equation 6] as long as $D\leq{D}_{\max}=0.15$. We attribute this behavior to the limitations in the domain of the Lagrangian multiplier $s_{P}$, as discussed in Theorem 8. We address this issue in detail in the following section.
@@ V-C On the convergence under TV perception constraint
As observed in Example 2, in the case of TV distance the RAM algorithm is not able to converge to proper solutions in the region $\Omega=\{(D,P)\in\mathbb{R}_{+}^{2}:D\geq p=0.15\}$. We argue that the problem stems from the values of the Lagrangian multipliers $s=(s_{D},s_{P})$ associated with $\Omega$ requires values of $s_{P}$ that do not guarantee the convergence of the algorithm, as reported in Theorem 8.
@@
$\mathcal{L}_{D_{f}}(D,P)\triangleq\{Q_{Y|X}\in\mathcal{Q}:\mathbb{E}_{Q_{Y|X}}\left[d(x,y)\right]\leq D\land D_{f}(p_{X}||q_{Y})\leq P\}.$
@@
$R_{D_{f}}(D,P)=\min_{Q_{Y|X}\in\mathcal{L}_{D_{f}}(D,P)}I(p_{X},Q_{Y|X})\quad\text{and}\quad R_{D_{g}}(D,P)=\min_{Q_{Y|X}\in\mathcal{L}_{D_{g}}(D,P)}I(p_{X},Q_{Y|X})$
@@
The results from Lemmas 7 and 8 guarantee that, $\forall n\in\mathbb{N}$, the RDP problem defined for the perception metric $D_{f_{n}}(\cdot||\cdot)$ act as lower-bound for the RDP problem defined with $TV(\cdot||\cdot)$ perception metric. Furthermore, since $D_{f_{n}}(\cdot||\cdot)$ is a smooth function, we can apply NAM scheme to compute the associate RDPF. Fig. 3 shows the RDPFs for a source $X\sim Ber(0.15)$ under Hamming distortion and perception measure $D_{f_{n}}$, for $n=\{1,10,100\}$. As expected, for increasing $n$, the $D_{f_{n}}$ perception metric becomes a better approximation of the $TV(\cdot||\cdot)$ perception.
@@
Figure 3: Comparison between the RDPF under $TV$ perception, computed with RAM scheme, and the RDPF under $D_{f_{n}}$, computed with NAM scheme, for $n$={1,10, 100}.
@@ VI Conclusion and Future Directions
In this paper, we studied the computation of the rate-distortion-perception function for discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint that belongs to the family of $f$-divergences. We showed that the introduction of the perception constraint induces a non-trivial generalization of the classical optimality conditions in terms of reconstruction distribution and transition matrix $(h^{*},Q_{Y|X}^{*})$, resulting in the OAM scheme. The presence of implicit equations in the formulation of the OAM scheme, implying the impossibility of direct algorithmic implementation, led to the formulation of alternative minimization schemes, NAM and RAM, each with its applicability conditions. Furthermore, we analyzed the asymptotic complexity of the proposed schemes and provided numerical results to validate our theoretical findings.
As an ongoing research, we plan to extend the derived optimality conditions to general measure spaces, as well as to study the behavior of the proposed algorithm when applied to continuous sources.
@@ Appendix A Proof of Lemma 1
The double minimization follows immediately from properties of the KL divergence, reported in [41, Theorem 5.2.6], which allows to rewrite (1) as (4). Moreover, the characterization (5) of the minimizer $h$ follows from the same properties.
On the other hand, the minimizing $Q_{Y|X}$ can be characterized by applying the Karush-Kuhn-Tucker (KKT) conditions [44] on the dual formulation of (1). The Lagrangian associated with the primal problem has form:
$\begin{split}L(Q_{Y|X},s,\lambda,\mu)&=D_{KL}\left(p_{X}Q_{Y|X}||p_{X}h\right)+s_{D}\left(\mathbb{E}\left[d(x,y)\right]-D\right)+s_{P}\left(D_{f}(p_{X}||q_{Y})-P\right)\\ &\quad+\sum_{x\in\mathcal{X}}\lambda_{x}\left(1-\sum_{y\in\mathcal{X}}Q_{Y|X}(y|x)\right)-\sum_{(x,y)\in\mathcal{X}^{2}}\mu_{x,y}Q_{Y|X}(y|x)\end{split}$
where the last two sets of constraints refer to the positivity and normalization constraints on $Q_{Y|X}$. Differentiating $L(Q_{Y|X},s,\lambda,\mu)$ with respect to the primal variables $Q_{Y|X}(y|x),\forall(x,y)\in\mathcal{X}^{2}$, we obtain:
$\begin{split}\frac{\partial L(Q_{Y|X},s,\lambda,\mu)}{\partial Q_{Y|X}(y|x)}=\frac{\partial D_{KL}(p_{X}Q_{Y|X}||p_{X}h)}{\partial Q_{Y|X}(y|x)}+s_{D}\frac{\partial E[d(X,Y)]}{\partial Q_{Y|X}(y|x)}+s_{P}\frac{\partial D_{f}(p_{X}||q_{Y})}{\partial Q_{Y|X}(y|x)}+\lambda_{x}-\mu_{x,y}\end{split}$
@@
$\begin{cases}\begin{split}\frac{\partial L(Q_{Y|X},s,\lambda,\mu)}{\partial Q_{Y|X}(y|x)}=0\\ \mu(x,y)Q_{Y|X}(y|x)=0\\ \end{split}\end{cases}$ (22)

### v2

@@ IV-A Asymptotic Convergence Rate Analysis
We note that the iteration structure in Theorem 2, i.e., $h^{(n+1)}$ = $q_{Y}[h^{(n)}]$ given the current iteration $n$, can be represented as an implicit vector function $S:\mathbb{R}^{|\mathcal{X}|}\to\mathbb{R}^{|\mathcal{X}|}$ with $S[h](i)=h(i)\cdot c[h,S[h]](i)$, such that $h^{(n+1)}=S[h^{(n)}]$. The results of Lemma 2 characterize a distribution $h^{*}$ that achieves the RDPF as a fixed point of $S(h)$, i.e., $h^{*}=S[h^{*}]$, since $c[h^{*},S[h^{*}]](i)=1,~i=1\ldots,|\mathcal{X}|$. Under these observations, we can analyze the convergence rate of the OAM scheme following similar steps as in [44].
The first-order Taylor expansion of $S[h]$ around a fixed point $h^{*}$ is defined as
[…]
@@
(Jacobian form) The Jacobian $J(h)$ evaluated at the fixed point $h^{*}$ is given as
@@
where $M^{*}=M[h^{*},h^{*}]$ and $\Gamma^{*}=\Gamma[h^{*},h^{*}]$ as defined in (14) and (15), respectively.
@@
Let $\{\lambda_{i}\}_{i\in[1:|\mathcal{X}|]}$ be the set of eigenvalues of $M^{*}=M[h^{*},h^{*}]$. Given a distortion function $d:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{0}^{+}$ that induces a full-rank matrix $D=[e^{-s_{D}d(i,j)}]_{(i,j)\in\mathcal{X}^{2}}$, then $\lambda_{i}>0$$\forall i\in[1:|\mathcal{X}|]$, i.e., $M^{*}$ has only positive eigenvalues.
@@
$0\leq\{\theta_{i}\}_{i\in[1:|\mathcal{X}|]}<1.$
[…]
@@
The convergence rate of the NAM scheme follows directly from the OAM scheme analysis, given the close relation between the two schemes. Since the only difference lies in the introduction of Newton’s root-finding method for the estimation of the optimal iteration step, the NAM scheme exhibits the same convergence rate in terms of the number of iterations as the OAM scheme, i.e., an exponential convergence $\mathcal{O}(e^{-n})$ under the assumptions of Lemma 5. However, the added complexity from the application of Newton’s method at each iteration increases the overall iteration complexity, due to the at least linear convergence rate $\mathcal{O}(\frac{1}{m})$ of the root approximant. Therefore, the total complexity is approximately $\mathcal{O}(\frac{e^{-n}}{m})$, where $n$ and $m$ depend on the error tolerances $\epsilon$ and $\delta$ given as input in Algo. 1.
@@
where $M$ and $\Gamma$ are given by (14) and (15), respectively. Unlike Theorem 2, where the structure of (21) bounds its own eigenvalues, in this case, we need to bound the Lagrangian multiplier $s_{P}$, hence matrix $\Gamma$, to guarantee exponential convergence of the algorithm. This is proved in the following theorem.
@@
For a given $s_{D}\geq{0}$, let $I_{s_{P}}=[0,s_{P,\max}]$ be the domain of $s_{P}$, $\{\theta_{a,i}\}_{i\in\mathcal{X}}$ the set of eigenvalues of $\hat{J}(h^{*})$ and $\theta_{\max}$ the maximum eigenvalue of $J(h^{*})$ in (21). Define the set $I_{s_{P}}^{\epsilon}=[0,s_{P,\max}-\epsilon]$ for $0<\epsilon<s_{P,\max}$. Then, there exists an $\epsilon^{\prime}$ such that if $s_{P}\in I_{s_{P}}^{\epsilon^{\prime}}$ then $0\leq\{\theta_{a,i}\}_{i\in\mathcal{X}}<1$.
@@ V Numerical Results
In this section, we validate our theoretical findings through simulations. In particular, we consider the computation of the RDPF under the NAM and RAM schemes, using Algorithms 1 and 2, respectively.
@@ V-A RDPF Computation - NAM scheme
Suppose that ${\mathcal{X}}=\{0,1\}$ with $p_{X}\sim Ber(0.15)$, and let $d(\cdot,\cdot)=d_{H}(\cdot,\cdot)$ with the perception constraint chosen as one of the following: (a) $D_{f}(\cdot||\cdot)=D_{JS}(\cdot||\cdot)$, (b) $D_{f}(\cdot||\cdot)=D_{KL}(\cdot||\cdot)$, (c) $D_{f}(\cdot||\cdot)=D_{\chi^{2}}(\cdot||\cdot)$, (d) $D_{f}(\cdot||\cdot)=D_{\alpha=-1}(\cdot||\cdot)$, (e) $D_{f}(\cdot||\cdot)=D_{\alpha=\frac{1}{2}}(\cdot||\cdot)$. In Fig. 1, we present the $R(D,P)$ estimates obtained using Algorithm 1 for each divergence metric.
@@
(e) $D_{\alpha=\frac{1}{2}}$
Figure 1: $R(D,P)$ for a Bernoulli source under Hamming distortion and various perception constraints.
We observe that all the computed RDPFs share similarities in the structure of the operating regions on the $(D,P)$ plane. Referring to the plot of Fig. 1a, we distinguish three cases:
  • Case I, where the perception constraint $P$ is not met with equality. In this region, the RDPF is equal to the RDF.
@@
The results show that Algorithm 1 can completely cover the Case III region. The boundary between Cases I-III (black line) representing the RDF is obtained by setting the Lagrangian multipliers $s_{P}=0$. Furthermore, the Case I region can be obtained by extension of the relative boundaries, since all the $(D,P)$ points share the solution given by the RD problem.
@@ V-B RDPF Computation - RAM scheme
Suppose that ${\mathcal{X}}=\hat{\cal X}=\{0,1\}$ with $p_{X}\sim Ber(0.15)$, and let $d(\cdot,\cdot)=d_{H}(\cdot,\cdot)$ with perception constraint chosen to be either (a) $D_{f}(p_{X}||h)=D_{JS}(\cdot||\cdot)$, (b) $D_{f}(\cdot||\cdot)=D_{KL}(\cdot||\cdot)$, (c) $D_{f}(\cdot||\cdot)=D_{\chi^{2}}(\cdot||\cdot)$, (d) $D_{f}(\cdot||\cdot)=D_{\alpha=-1}(\cdot||\cdot)$, (e) $D_{f}(\cdot||\cdot)=TV(\cdot||\cdot)$. In Fig. 2, we present the $R(D,P)$ estimates obtained using Algorithm 2 for each divergence metric.
@@
(e) $TV(\cdot||\cdot)$
Figure 2: $R(D,P)$ for a Bernoulli source under Hamming distortion and various perception constraints.
Similar remarks to those in Example 1 apply to the operating regions of the computed RDPFs. In addition, in Fig. 2e we compare the theoretical results of [3, Equation 6] with the numerical results obtained using Algorithm 2. This case is of particular interest due to the non-differentiability of the $TV$ distance. We observe that Algorithm 2 achieves exactly the theoretical solution of [3, Equation 6] as long as $D\leq{D}_{\max}=0.15$. We attribute this behavior to limitations in the domain of the Lagrangian multiplier $s_{P}$, as discussed in Theorem 8. We address this issue in detail in the following section.
@@ V-C On the convergence under TV perception constraint
As observed in Example 2, in the case of TV distance, the RAM algorithm is not able to converge to proper solutions in the region $\Omega=\{(D,P)\in\mathbb{R}_{+}^{2}:D\geq p=0.15\}$. We argue that the problem stems from the values of the Lagrangian multipliers $s=(s_{D},s_{P})$ associated with $\Omega$ require values of $s_{P}$ that do not guarantee the convergence of the algorithm, as reported in Theorem 8.
@@
$\mathcal{L}_{D_{f}}(D,P)\triangleq\{Q_{Y|X}:\mathbb{E}_{Q_{Y|X}}\left[d(X,Y)\right]\leq D,D_{f}(p_{X}||q_{Y})\leq P\}.$
@@
$R_{D_{f}}(D,P)=\min_{Q_{Y|X}\in\mathcal{L}_{D_{f}}(D,P)}I(p_{X},Q_{Y|X})$
$R_{D_{g}}(D,P)=\min_{Q_{Y|X}\in\mathcal{L}_{D_{g}}(D,P)}I(p_{X},Q_{Y|X})$
@@
The results from Lemmas 7 and 8 guarantee that, $\forall n\in\mathbb{N}$, the RDP problem defined for the perception metric $D_{f_{n}}(\cdot||\cdot)$ acts as a lower bound for the RDP problem defined with $TV(\cdot||\cdot)$ perception metric. Furthermore, since $D_{f_{n}}(\cdot||\cdot)$ is a smooth function, we can apply the NAM scheme to compute the associated RDPF. Fig. 3 shows the RDPFs for a source $X\sim Ber(0.15)$ under Hamming distortion and perception measure $D_{f_{n}}$, for $n\in\{1,10,100\}$. As expected, for increasing $n$, $D_{f_{n}}$ provides a progressively better approximation of the $TV(\cdot||\cdot)$ metric.
@@
Figure 3: Comparison between the RDPF under TVTV perception, computed with the RAM scheme, and the RDPF under $D_{f_{n}}$, computed with the NAM scheme, for $n\in\{1,10,100\}$.
@@ VI Conclusion and Future Directions
In this paper, we have studied the computation of the rate-distortion-perception function for discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint from the family of $f$-divergences. We showed that the introduction of the perception constraint induces a non-trivial generalization of the classical optimality conditions for the reconstruction distribution and the transition matrix $(h^{*},Q_{Y|X}^{*})$, resulting in the OAM scheme. Since the OAM scheme involves implicit equations that preclude direct algorithmic implementation, we introduced alternative minimization schemes, NAM and RAM, each with its own applicability conditions. Furthermore, we analyzed the asymptotic complexity of the proposed schemes and provided numerical results to validate our theoretical findings.
As ongoing research, we plan to extend the derived optimality conditions to general measure spaces, as well as to study the behavior of the proposed algorithms when applied to continuous sources.
@@ Appendix A Proof of Lemma 1
The double minimization follows immediately from properties of the KL divergence, reported in [43, Theorem 5.2.6], which allows to rewrite (1) as (4). Moreover, the characterization (5) of the minimizer $h$ follows from the same properties.
On the other hand, we remark that (1) is a convex program in the variable $Q_{Y|X}$ for a given $p_{X}$, and respects Slater’s condition, since it is easy to show that $Q_{Y|X}(y|x)=\delta_{y=x}$ is an interior point of the constraint set, as highlighted in Remark 2. Therefore, the minimizer $Q_{Y|X}^{*}$ can be characterized by applying the Karush-Kuhn-Tucker (KKT) conditions [46] on the dual formulation of (1).
The Lagrangian associated with the primal problem has the form:
$L(Q_{Y|X},s,\lambda,\mu)$
$=D_{KL}\left(p_{X}Q_{Y|X}||p_{X}h\right)+s_{D}\left(\mathbb{E}\left[d(X,Y)\right]-D\right)$
$\quad+s_{P}\left(D_{f}(p_{X}||q_{Y})-P\right)+\sum_{x\in\mathcal{X}}\lambda_{x}\left(1-\sum_{y\in\mathcal{X}}Q_{Y|X}(y|x)\right)$
$\quad-\sum_{(x,y)\in\mathcal{X}^{2}}\mu_{x,y}Q_{Y|X}(y|x)$
where the last two sets of constraints refer to the positivity and normalization constraints on $Q_{Y|X}$. Differentiating $L(Q_{Y|X},s,\lambda,\mu)$ with respect to the primal variables $Q_{Y|X}(y|x),\forall(x,y)\in\mathcal{X}^{2}$, we obtain
$\begin{split}&\frac{\partial L(Q_{Y|X},s,\lambda,\mu)}{\partial Q_{Y|X}(y|x)}\\ &~~~~~~~~~=\frac{\partial D_{KL}(p_{X}Q_{Y|X}||p_{X}h)}{\partial Q_{Y|X}(y|x)}+s_{D}\frac{\partial E[d(X,Y)]}{\partial Q_{Y|X}(y|x)}\\ &~~~~~~~~~~~~~+s_{P}\frac{\partial D_{f}(p_{X}||q_{Y})}{\partial Q_{Y|X}(y|x)}+\lambda_{x}-\mu_{x,y}\end{split}$
@@
[…]